OpenStudy (midhun.madhu1987):

999 1/7 + 999 2/7 + 999 3/7 + 999 4/7 + 999 5/7 + 999 6/7 What is the sum of them?

OpenStudy (midhun.madhu1987):

This is Mixed fraction...

OpenStudy (anonymous):

\[999+\frac{1}{7}+\cdots+999+\frac{6}{7}=6(999)+\frac{21}{7}=3(2\times999+1)\]

OpenStudy (gorv):

999+1/7............+999+6/7+.......

OpenStudy (midhun.madhu1987):

Got it... Thanks :) all